OpenStudy (nali):

solve the equation 3 tan x - cot x = 0

OpenStudy (anonymous):

Let tan x be equal to u and then solve for u

OpenStudy (anonymous):

Remember \[\cot x = 1/tanx\]

OpenStudy (anonymous):

3 tan x = cot x 3 tan^2(x) = 1 tan^2(x) = 1/3 \[\tan x = \sqrt{3}/3\] x = 30

OpenStudy (anonymous):

so you have \[3u - 1/u = 0\]

OpenStudy (nali):

what do u mean by u u and x are the same thing

OpenStudy (anonymous):

Look at the comment above yours. I wrote it down for you.

OpenStudy (nali):

ok one

OpenStudy (nali):

question how did u get tan^2 @jlee7x2

OpenStudy (anonymous):

No, your answer would not be one.

OpenStudy (anonymous):

3 tan x = 1/tan x 3 tan^2(x) = 1

OpenStudy (nali):

yes i know that it was a typo

OpenStudy (anonymous):

b/c cot x = 1/tan x as @beeqay mentioned

OpenStudy (anonymous):

\[3u- 1/u =0\] This is equal to \[3u^2 -1= 0\] Now solve for u

OpenStudy (nali):

how did u get tan^2 from tan x - 1/tan x

OpenStudy (anonymous):

@Nali If you plug in \[tanx \]back into \[u\], then you have \[3\tan^2x -1 = 0\]

OpenStudy (anonymous):

But @jlee7x2 just multiplied everything by\[ tanx\] to get rid of \[1/tanx\] You see \[tanx- 1/tanx=0\] is the same as \[\tan^2x -1 = 0\]

OpenStudy (nali):

yes i got it thanks to all

OpenStudy (anonymous):

No problem.